# Graph Neural Network Enhanced Retrieval for Question Answering of LLMs

### v1

@@ Appendix F Prompts for Experiments
F.1 Prompt template for Direct.
@@
Prompt template for Direct Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words.
<Question>
@@
F.2 Prompt templates for Dense, IRCoT, and KGP
Dense, IRCoT, and KGP employ different retrieval methods but the same prompt templates for QA. The prompt template is shown as follows:
Prompt template for Dense / IRCoT / KGP Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words.
Context: <Context>
Question: <Question>
@@ F.2 Prompt templates for Dense, IRCoT, and KGP
F.3 Prompt template for SelfAsk
@@
Prompt template for SelfAsk Instruction: Your goal is to answer the question step by step following procedures of examples. If no follow up questions are necessary, answer the question in five words directly in format: So the final answer is: xxx.
Here are the examples of how to answer the questions:
<Examples>
Context (Optional): <Context>
Question:
<Question>
@@
F.4 Prompt template for ITER-RETGEN
@@
Prompt template for ITER-RETGEN Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words.
Here are the examples of how to answer the questions:
<Examples>
Context: <Context>
Question:
<Question>
Let’s think step by step.
F.5 Prompt template for RGNN-Ret
@@
Prompt template for generating next-step subquestions Instruction: Your goal is to ask the next step question logically.
Here are the examples of how to answer the questions:
<Examples>
Question:
<Question>
Prompt template for generating subanswers Instruction: Your goal is to answer the next step question. I will provide you some wikipedia snippets, and you need to answer the next step question by considering the wikipedia snippets.
Context:
<Context>
Question:
<Question>
@@
Prompt template of self-critique Instruction: You are a wikipedia QA expert. Your goal is to critique whether the intermediate answers are enough to generate the final answer to the question.First analyze if the intermediate answers is enough to generate the final answer step by step logically. Then, if it is enough, output ’Critique: yes’. If not, output ’Critique: no’.
Question:
<Question>
"Analyze if the intermediate answers is enough to generate the final answer to the question step by step logically. Then, if it is enough, output ’Critique: yes’. If not, output ’Critique: no’.
@@
Demonstrations of RGNN-Ret for asking next-step subquestions on MuSiQue dataset Question: Who lived longer, Muhammad Ali or Alan Turing?
Next step: How old was Muhammad Ali when he died?
Intermediate answer: Muhammad Ali was 74 years old when he died.
Next step: How old was Alan Turing when he died?
Question: When was the founder of craigslist born?
Next step: Who was the founder of craigslist?
Intermediate answer: Craigslist was founded by Craig Newmark.
Next step: When was Craig Newmark born?
Question: Who was the maternal grandfather of George Washington?
Next step: Who was the mother of George Washington?
Intermediate answer: The mother of George Washington was Mary Ball Washington.
Next step: Who was the father of Mary Ball Washington?
Question: Are both the directors of Jaws and Casino Royale from the same country?
Next step: Who is the director of Jaws?
Intermediate Answer: The director of Jaws is Steven Spielberg.
Next step: Where is Steven Spielberg from?
Intermediate Answer: The United States.
Next step: Who is the director of Casino Royale?
Intermediate Answer: The director of Casino Royale is Martin Campbell.
Next step: Where is Martin Campbell from?
Demonstrations of RGNN-Ret for asking next-step subquestions on IIRC dataset Question: Who lived longer, Muhammad Ali or Alan Turing?
Next step: How old was Muhammad Ali when he died?
Intermediate answer: Muhammad Ali was 74 years old when he died.
Next step: How old was Alan Turing when he died?
Question: When was the founder of craigslist born?
Next step: Who was the founder of craigslist?
Intermediate answer: Craigslist was founded by Craig Newmark.
Next step: When was Craig Newmark born?
Question: Was the city where Eva was born the capital of its country?
Next step: Where was Eva born?
Intermediate answer: Eva was born in Berlin, Germany.
Next step: Is Berlin the capital of Germany?
Question: Was Ryuji Yamakawa a good solo wrestler?
Next step: Who is Ryuji Yamakawa?
Intermediate answer: Ryuji Yamakawa is a Japanese professional wrestler.
Next step: Was Ryuji Yamakawa a good solo wrestler?
Demonstrations of RGNN-Ret for asking next-step subquestions on 2WikiMQA dataset Question: Who lived longer, Muhammad Ali or Alan Turing?
Next step: How long did Muhammad Ali live?
Intermediate answer: Muhammad Ali was 74 years old when he died.
Next step: How long did Alan Turing live?
Question: Who was the paternal grandfather of Princess Alexandrine Of Prussia (1842-1906)?
Next step: Who was the father of Princess Alexandrine Of Prussia (1842-1906)?
Intermediate answer: Princess Alexandrine Of Prussia (1842-1906) was the daughter of Prince Albert of Prussia.
Next step: Who was the father of Prince Albert of Prussia?
Question: Who is the mother of the composer of film 404 (Film)?
Next step: Who is the composer of film 404 (Film)?
Intermediate answer: The composer of film 404 (Film) is Ilayaraja.
Next step: Who is the mother of Ilayaraja?
Question: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality?
Next step: What is the nationality of the director of Across The Badlands?
Intermediate answer: The director of Across The Badlands is American.
Next step: What is the nationality of the director of A Gutter Magdalene?
Intermediate answer: The director of A Gutter Magdalene is American.
Next step: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality?
Question: What is the home stadium of the team that Asprey hit a hat trick against on 16 January 1961?
Next step: Which team did Asprey hit a hat trick against on 16 January 1961?
Intermediate answer: Asprey hit a hat trick against Charlton Athletic on 16 January 1961.
Next step: What is the home stadium of Charlton Athletic?
Appendix G Training data for RGNN
@@
Training data of MuSiQue for RGNN 1. Question: Why did Roncali leave the place of death of the creator of Malchiostro Annunciation?
Subquestions: Who is the creator of Malchiostro Annunciation? Where did Titian die? Why did Roncali leave Venice?
2. Question: Where did who argued that the country of citizenship of Victor Denisov had itself beome an imperialist power declare that he would intervene in the Korean conflict?
Subquestions: What is the country of citizenship of Victor Denisov? Who argued that Russia had itself become an imperialist power? Where did Mao Zedong declare that he would intervene in the Korean conflict?
3. Question: What military overran much of Erich Zakowski’s place of birth?
Subquestions: What is the place of birth of Erich Zakowski? What military overran East Prussia?
4. Question: How many people were in British Colonies where does the london broil cut come from ?
Subquestions: Where does the london broil cut come from? How many people were in North American?
5. Question: When was the country established that lies immediately north of the Persian Gulf and the region where the country containing Urim is located?
Subquestions: What is the region containing Urim? Where is the region that Iraq is located? What is the country that lies immediately north of the Persian Gulf and the Middle East? When was Iran established?
Training data of IIRC for RGNN 1. Question: Was the city where Eva was born the capital of its country?
Subquestions: Where was Eva born? Is Berlin the capital of its country?
2. Question: In what state did Galambos attend high school?
Subquestions: What high school did Galambos attend? In what state is Athens High School located?
3. Question: What was the previous name of the team that Feng started playing with in 1999?
Subquestions: What team did Feng start playing with in 1999? What was the previous name of Chongqing Longxin?
4. Question: How many years after the first Marvel Cinematic Universe film came out was Black Panther released?
Subquestions: When was the first Marvel Cinematic Universe film released? When was Black Panther released? How many years after Iron Man came out was Black Panther released?
5. Question: Who was the first draft pick the year Damarius Bilbo went undrafted?
Subquestions: What year did Damarius Bilbo go undrafted? Who was the first draft pick in 2006?
Training data of 2WikiMQA for RGNN 1. Question: Which film came out earlier, Subliminal Seduction or Australia Marches With Britain?
Subquestions: When did Subliminal Seduction come out? When did Australia Marches With Britain come out?
2. Question: Who is the father-in-law of Deuteria?
Subquestions: Who is the husband of Deuteria? Who is the father of Eusebio?
3. Question: Who is the mother of the composer of film 404 (Film)?
Subquestions: Who is the composer of film 404 (Film)? Who is the mother of Ilayaraja?
4. Question: Which country the composer of film Sergeant Hassan is from?
Subquestions: Who is the composer of film Sergeant Hassan? Which country is Tamer Karaoğlu from?
5. Question: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality?
Subquestions: What is the director of Across The Badlands? Where is Fred F. Sears from? What is the director of A Gutter Magdalene? Where is George Melford from?

### v2

@@ Appendix I Prompts for Experiments
I.1 Prompt template for Direct.
@@
Prompt template for Direct Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words. <Question>
@@
I.2 Prompt templates for dense retrievers
The prompt template for retrievers (bm25, DPR, and SBERT) is shown as follows:
Prompt template for retrievers Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words. Context: <Context> Question: <Question>
@@ I.2 Prompt templates for dense retrievers
I.3 Prompt templates for IRCoT and KGP
IRCoT and KGP employ different retrieval methods but the same prompt templates for QA. The prompt template is shown as follows:
Prompt template for IRCoT / KGP Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words. Context: <Context> Question: <Question>
where <Context> indicates the retrieved passages.
I.4 Prompt template for SelfAsk
@@
Prompt template for SelfAsk Instruction: Your goal is to answer the question step by step following procedures of examples. If no follow up questions are necessary, answer the question in five words directly in format: So the final answer is: xxx. Here are the examples of how to answer the questions: <Examples> Context (Optional): <Context> Question: <Question>
@@
I.5 Prompt template for ITER-RETGEN
@@
Prompt template for ITER-RETGEN Instruction: Given the following question, create a final answer to the question. Please answer less than 6 words. Here are the examples of how to answer the questions: <Examples> Context: <Context> Question: <Question> Let’s think step by step.
I.6 Prompt template for RGNN-Ret
@@
Prompt template for generating next-step subquestions Instruction: Your goal is to ask the next step question logically. Here are the examples of how to answer the questions: <Examples> Question: <Question>
Prompt template for generating subanswers Instruction: Your goal is to answer the next step question. I will provide you some wikipedia snippets, and you need to answer the next step question by considering the wikipedia snippets. Context: <Context> Question: <Question>
@@
Prompt template of self-critique Instruction: You are a wikipedia QA expert. Your goal is to critique whether the intermediate answers are enough to generate the final answer to the question.First analyze if the intermediate answers is enough to generate the final answer step by step logically. Then, if it is enough, output ’Critique: yes’. If not, output ’Critique: no’. Question: <Question> "Analyze if the intermediate answers is enough to generate the final answer to the question step by step logically. Then, if it is enough, output ’Critique: yes’. If not, output ’Critique: no’.
@@
Demonstrations of RGNN-Ret for asking next-step subquestions on MuSiQue dataset Question: Who lived longer, Muhammad Ali or Alan Turing? Next step: How old was Muhammad Ali when he died? Intermediate answer: Muhammad Ali was 74 years old when he died. Next step: How old was Alan Turing when he died? Question: When was the founder of craigslist born? Next step: Who was the founder of craigslist? Intermediate answer: Craigslist was founded by Craig Newmark. Next step: When was Craig Newmark born? Question: Who was the maternal grandfather of George Washington? Next step: Who was the mother of George Washington? Intermediate answer: The mother of George Washington was Mary Ball Washington. Next step: Who was the father of Mary Ball Washington? Question: Are both the directors of Jaws and Casino Royale from the same country? Next step: Who is the director of Jaws? Intermediate Answer: The director of Jaws is Steven Spielberg. Next step: Where is Steven Spielberg from? Intermediate Answer: The United States. Next step: Who is the director of Casino Royale? Intermediate Answer: The director of Casino Royale is Martin Campbell. Next step: Where is Martin Campbell from?
Demonstrations of RGNN-Ret for asking next-step subquestions on IIRC dataset Question: Who lived longer, Muhammad Ali or Alan Turing? Next step: How old was Muhammad Ali when he died? Intermediate answer: Muhammad Ali was 74 years old when he died. Next step: How old was Alan Turing when he died? Question: When was the founder of craigslist born? Next step: Who was the founder of craigslist? Intermediate answer: Craigslist was founded by Craig Newmark. Next step: When was Craig Newmark born? Question: Was the city where Eva was born the capital of its country? Next step: Where was Eva born? Intermediate answer: Eva was born in Berlin, Germany. Next step: Is Berlin the capital of Germany? Question: Was Ryuji Yamakawa a good solo wrestler? Next step: Who is Ryuji Yamakawa? Intermediate answer: Ryuji Yamakawa is a Japanese professional wrestler. Next step: Was Ryuji Yamakawa a good solo wrestler?
Demonstrations of RGNN-Ret for asking next-step subquestions on 2WikiMQA dataset Question: Who lived longer, Muhammad Ali or Alan Turing? Next step: How long did Muhammad Ali live? Intermediate answer: Muhammad Ali was 74 years old when he died. Next step: How long did Alan Turing live? Question: Who was the paternal grandfather of Princess Alexandrine Of Prussia (1842-1906)? Next step: Who was the father of Princess Alexandrine Of Prussia (1842-1906)? Intermediate answer: Princess Alexandrine Of Prussia (1842-1906) was the daughter of Prince Albert of Prussia. Next step: Who was the father of Prince Albert of Prussia? Question: Who is the mother of the composer of film 404 (Film)? Next step: Who is the composer of film 404 (Film)? Intermediate answer: The composer of film 404 (Film) is Ilayaraja. Next step: Who is the mother of Ilayaraja? Question: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality? Next step: What is the nationality of the director of Across The Badlands? Intermediate answer: The director of Across The Badlands is American. Next step: What is the nationality of the director of A Gutter Magdalene? Intermediate answer: The director of A Gutter Magdalene is American. Next step: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality? Question: What is the home stadium of the team that Asprey hit a hat trick against on 16 January 1961? Next step: Which team did Asprey hit a hat trick against on 16 January 1961? Intermediate answer: Asprey hit a hat trick against Charlton Athletic on 16 January 1961. Next step: What is the home stadium of Charlton Athletic?
Appendix J Training data for RGNN
@@
Training data of MuSiQue for RGNN 1. Question: Why did Roncali leave the place of death of the creator of Malchiostro Annunciation? Subquestions: Who is the creator of Malchiostro Annunciation? Where did Titian die? Why did Roncali leave Venice? 2. Question: Where did who argued that the country of citizenship of Victor Denisov had itself beome an imperialist power declare that he would intervene in the Korean conflict? Subquestions: What is the country of citizenship of Victor Denisov? Who argued that Russia had itself become an imperialist power? Where did Mao Zedong declare that he would intervene in the Korean conflict? 3. Question: What military overran much of Erich Zakowski’s place of birth? Subquestions: What is the place of birth of Erich Zakowski? What military overran East Prussia? 4. Question: How many people were in British Colonies where does the london broil cut come from ? Subquestions: Where does the london broil cut come from? How many people were in North American? 5. Question: When was the country established that lies immediately north of the Persian Gulf and the region where the country containing Urim is located? Subquestions: What is the region containing Urim? Where is the region that Iraq is located? What is the country that lies immediately north of the Persian Gulf and the Middle East? When was Iran established?
Training data of IIRC for RGNN 1. Question: Was the city where Eva was born the capital of its country? Subquestions: Where was Eva born? Is Berlin the capital of its country? 2. Question: In what state did Galambos attend high school? Subquestions: What high school did Galambos attend? In what state is Athens High School located? 3. Question: What was the previous name of the team that Feng started playing with in 1999? Subquestions: What team did Feng start playing with in 1999? What was the previous name of Chongqing Longxin? 4. Question: How many years after the first Marvel Cinematic Universe film came out was Black Panther released? Subquestions: When was the first Marvel Cinematic Universe film released? When was Black Panther released? How many years after Iron Man came out was Black Panther released? 5. Question: Who was the first draft pick the year Damarius Bilbo went undrafted? Subquestions: What year did Damarius Bilbo go undrafted? Who was the first draft pick in 2006?
Training data of 2WikiMQA for RGNN 1. Question: Which film came out earlier, Subliminal Seduction or Australia Marches With Britain? Subquestions: When did Subliminal Seduction come out? When did Australia Marches With Britain come out? 2. Question: Who is the father-in-law of Deuteria? Subquestions: Who is the husband of Deuteria? Who is the father of Eusebio? 3. Question: Who is the mother of the composer of film 404 (Film)? Subquestions: Who is the composer of film 404 (Film)? Who is the mother of Ilayaraja? 4. Question: Which country the composer of film Sergeant Hassan is from? Subquestions: Who is the composer of film Sergeant Hassan? Which country is Tamer Karaoğlu from? 5. Question: Do both films Across The Badlands and A Gutter Magdalene have the directors that share the same nationality? Subquestions: What is the director of Across The Badlands? Where is Fred F. Sears from? What is the director of A Gutter Magdalene? Where is George Melford from?